# MT-RewardTree: A Comprehensive Framework for Advancing LLM-Based Machine Translation via Reward Modeling

### v1

@@ Abstract
Process reward models (PRMs) have shown success in complex reasoning tasks for large language models (LLMs). However, their application to machine translation (MT) remains underexplored due to the lack of systematic methodologies and evaluation benchmarks. To address this gap, we introduce MT-RewardTree, a comprehensive framework for constructing, evaluating, and deploying process reward models in MT. Unlike traditional vanilla preference pair construction, we propose a novel method for automatically generating token-level preference pairs using approximate Monte Carlo Tree Search (MCTS), which mitigates the prohibitive cost of human annotation for fine-grained steps. Then, we establish the first MT-specific reward model benchmark and provide a systematic comparison of different reward modeling architectures, revealing that token-level supervision effectively captures fine-grained preferences. Experimental results demonstrate that our MT-PRM-Qwen-2.5-3B achieves state-of-the-art performance in both token-level and sequence-level evaluation given the same input prefix. Furthermore, we showcase practical applications where PRMs enable test-time alignment for LLMs without additional alignment training and significantly improve performance in hypothesis ensembling. Our work provides valuable insights into the role of reward models in MT research. Our code and data are released in https://sabijun.github.io/MT_RewardTreePage.
@@
Zhaopeng Feng1♠ Jiahan Ren1♠  Jiayuan Su1♠  Jiamei Zheng1 Zhihang Tang1  Hongwei Wang1†  Zuozhu Liu1† 1Zhejiang University {zhaopeng.23,jiahan.24,jiayuan.23,jiamei.24,hongweiwang,zuozhuliu}@intl.zju.edu.cn
@@ 1 Introduction
Machine translation (MT) naturally aligns with token-level MDP frameworks, as each translation decision corresponds directly to token generation. Ramos et al. (2024) pioneered the use of the automatic metric xCOMET (Guerreiro et al., 2024) as a reward model to generate dynamic token-level rewards for RL training. Additionally, Zhao et al. (2024) and Wang et al. (2024b) demonstrated that scaling test-time compute can enhance LLM-based translation quality. However, these works do not fully address reward modeling, and there is still a lack of systematic methodologies for constructing and evaluating PRMs in MT, which has hindered progress relative to advancements in general-domain LLMs.
@@
Developing effective PRMs is challenging. Although Lightman et al. (2024) demonstrate that process supervision with human annotators improves PRM performance in mathematical tasks, this methodology requires domain-expert annotators, resulting in prohibitive costs and practical limitations for translation tasks. Recently, some studies suggest that a PRM can be automatically learned during Direct Preference Optimization (DPO) training (Rafailov et al., 2024b, a; Yuan et al., 2024). However, existing vanilla preference pair datasets provide only sequence-to-sequence preference data, rather than token-level preferences, which raises concerns about their applicability for token-level alignment. Additionally, evaluating PRMs remains a significant challenge. In mathematical tasks, evaluation is often done using a Best-of-N (BoN) sampling strategy—selecting the highest-scored response from N candidates based on a PRM (Lightman et al., 2024; Wang et al., 2024c; Luo et al., 2024)—or by having the PRM identify errors or verify correctness in the steps (Zheng et al., 2024; Zhang et al., 2025). Since each step in mathematics has a deterministic answer, these methods do not directly translate to PRM evaluation in MT.
@@
In this paper, we introduce MT-RewardTree, a comprehensive framework for constructing, evaluating, and deploying PRMs in machine translation. We propose an approximate Monte Carlo Tree Search (MCTS) method (Kocsis and Szepesvári, 2006; Silver et al., 2016) to generate the token-level preference pair dataset. This dataset is then split into a training set for reward model development and a benchmark for reward evaluation. We provide a systematic comparison of different reward modeling methods and test on both token-level and sequence-level performance. Furthermore, we demonstrate two practical applications of PRMs, including test-time alignment and hypothesis ensembling, offering valuable insights for future MT research. Our main contributions are as follows:
  • We construct the first reward model benchmark in MT and provide the first systematic comparison of reward modeling approaches for MT. Experimental results show that our MT-PRMs achieve competitive performance at both the token-level and sequence-level.
@@
  • Our token-level preference pairs, generated through an approximate MCTS method, significantly outperform vanilla preference pairs in process reward model training.
  • We demonstrate that our MT-PRMs can facilitate test-time alignment for LLM-based MT without the need for additional alignment training. They can also be integrated for hypothesis ensembling to provide better translations, delivering performance parity with both MBR decoding and commercial LLM selection.
@@ 2.1 Token-level Markov Decision Process
where $(x_{0},\dots,x_{L})$ represents the input prompt and $(y_{0},\dots,y_{t-1})$ is the sequence of generated tokens until time step $t-1$. The state transition function $f$ is deterministic and updates the state by concatenating the newly generated token:
@@ 2.2 Reward Modeling in RLHF
Although token-level reward modeling offers finer-grained feedback, obtaining effective PRMs is more challenging to obtain and deploy compared to ORMs (Lightman et al., 2024; Cao et al., 2024).
@@ 3.1 Constructing Token-level Preference Pairs
Drawing inspiration from MCTS, we propose a token-centric approach that quantifies token quality based on its potential to contribute to higher-quality translations. This method aligns with Monte Carlo-based PRMs construction techniques in mathematics, where step-wise quality is determined by its incremental contribution to deriving correct answers (Wang et al., 2024c; Guan et al., 2025).
@@
Model Sequence-level Token-level
Prefixed Arbitrary Prefixed
EN→XX XX→EN Avg. EN→XX XX→EN Avg. EN→XX XX→EN Avg.
@@
Skywork-Reward-LLaMA-3.1-8B 0.857 0.773 0.815 0.768 0.743 0.756 - - -
MT-Ranker-base 0.785 0.787 0.786 0.720 0.712 0.716 - - -
MT-Ranker-large 0.847 0.873 0.860 0.750 0.805 0.778 - - -
[…]
MT-PRM-LLaMA-3.2-3B 0.777 0.775 0.776 0.675 0.660 0.668 0.542 0.615 0.578
MT-PRM-Qwen-2.5-3B 0.867 0.858 0.863 0.622 0.633 0.628 0.637 0.685 0.660
Table 1: Accuracy results on RM-RewardBench. Skywork-Reward-LLaMA-3.1-8B is an advanced ORM for general domains, while MT-Ranker represents the SoTA non-metric reference-free translation quality estimation model.
Preference Pair Type Training Strategy Avg.
Token-level DPO 0.660
Vanilla DPO 0.574
Token-level KTO 0.644
Vanilla KTO 0.562
Table 2: Ablation study on the effect of training preference data and implicit reward training objectives. The backbone model is Qwen-2.5-3B-Instruct.
@@ 3.2 Implicit Process Reward Modeling
4 Implementation Details
Datasets. We explore four languages—English (EN), German (DE), Chinese (ZH), and Russian (RU)—and six translation directions: EN→XX and XX→EN. Our raw corpus consists of test sets from WMT17 to WMT20, supplemented with development and test sets from the Flores (Costa-jussà et al., 2022). We use the TowerInstruct-7B-v0.2111https://huggingface.co/Unbabel/TowerInstruct-7B-v0.2 model with a temperature of 0.95 and apply the MCTS-based approach described earlier. During the Simulation step, we sample three candidate hypotheses for each node.
The token-level preference pair dataset is divided into two subsets: 200 pairs per translation direction are sampled as the token-level MT-RewardBench (Prefixed), which also serves as the sequence-level MT-RewardBench (Prefixed). The remaining pairs are assigned as token-level training data. Additionally, we sample two translations for each source sentence in the token-level datasets, ensuring they do not strictly share the same prefixed text (namely the vanilla preference pair). These pairs form the sequence-level arbitrary preference pair set. Detailed statistics can be found in Table 5.
Training Details. We take LLaMA-3.2-3B-Instruct [^2] https://huggingface.co/meta-LLaMA/LLaMA-3.2-3B-Instruct and Qwen-2.5-3B-Instruct [^3] https://huggingface.co/Qwen/Qwen2.5-3B-Instruct as the backbone models for training. For the DPO training, the higher-scored sentence is designated as the chosen response, while the lower-scored sentence is labeled as the rejected response. For the KTO training, the higher-scored sentence is treated as the positive sample, and the lower-scored sentence as the negative sample. We set $\beta$ as 0.1.
@@ 4 Implementation Details
5 Evaluation Results
@@
Token-level Performance. From Table 3.1, we can observe that our MT-PRM-LLaMA-3.2-3B and MT-PRM-Qwen-2.5-3B models achieved accuracies of 0.578 and 0.66 respectively on the token-level MT-RewardBench. As shown in Table 2, we systematically compare models trained with vanilla sequence-level preference pairs versus our token-level preference pairs, while evaluating both DPO and KTO training objectives. The results demonstrate that token-level preference pairs significantly improve discrimination accuracy: implicit PRMs trained with token-level preference pairs outperform vanilla sequence-level baselines by +8.6% (DPO) and +11.5% (KTO). This performance gap highlights the critical advantage of token-level preference pairs in helping capture fine-grained translation quality distinctions.
@@
Sequence-level Performance. We also convert our PRMs to sequence-level scoring through weighted DPO rewards (as shown in Eq. 11). We can observe that our MT-PRM-Qwen-2.5-3B achieves the highest performance among all models in the Prefixed set, with an average score of 0.863, outperforming both Skywork-Reward-LLaMA-3.1-8B  [^4] https://huggingface.co/Skywork/Skywork-Reward-Llama-3.1-8B and the MT-Ranker (Moosa et al., 2024) variants. This demonstrates the effectiveness of our token-level supervision framework even when adapted to sequence-level scoring. However, in the Arbitrary set, we observe a relative drop in performance (average accuracy between 0.628 and 0.668). We hypothesize this drop stems from the inclusion of noisy partial translations in the Arbitrary set, which disrupts the causal dependency structure crucial for token-level rewards, and the limited compositional reasoning capabilities of current PRM architectures when handling fragmented inputs, a challenge well-documented in partial sequence evaluation (Wang et al., 2024a; Yuan et al., 2024).
Strategy Model ZH-EN EN-ZH Avg.
XCOMET COMETKiwi XCOMET COMETKiwi
MBR BLEU 0.8243 0.7842 0.8106 0.7768 0.7990
BERTScore 0.8277 0.7874 0.8227 0.7837 0.8054
COMET 0.8371 0.7932 0.8334 0.7949 0.8147
[…]
MT-PRM-Qwen2.5-3B 0.8250 0.7866 0.8153 0.7841 0.8028
MT-PRM-LLaMA3.2-3B 0.8282 0.7908 0.8225 0.7918 0.8083
Table 3: Automatic evaluation metrics for different ensembling strategies across WMT 23 ZH-EN and EN-ZH.
Source {CJK*}UTF8gkai 换油的师傅说油品清亮,确实是好油。
[…]
Table 4: Case study of test-time alignment.
6 Practical Insights

### v2

@@ Abstract
Process reward models (PRMs) have shown success in complex reasoning tasks for large language models (LLMs). However, their application to machine translation (MT) remains underexplored due to the lack of systematic methodologies and evaluation benchmarks. To address this gap, we introduce MT-RewardTree, a comprehensive framework for constructing, evaluating, and deploying process reward models in MT. Unlike traditional vanilla preference pair construction, we propose a novel method for automatically generating token-level preference pairs using approximate Monte Carlo Tree Search (MCTS), which mitigates the prohibitive cost of human annotation for fine-grained steps. Then, we establish the first MT-specific reward model benchmark and provide a systematic comparison of different reward modeling architectures, revealing that token-level supervision effectively captures fine-grained preferences. Experimental results demonstrate that our MT-PRM-Qwen-2.5-3B achieves state-of-the-art performance in both token-level and sequence-level evaluation given the same input prefix. Furthermore, we showcase practical applications where MT-PRMs successfully identify token-level translation differences and enable test-time alignment for LLMs without additional alignment training. Our work provides valuable insights into the role of reward models in MT research. Our code and data are released in https://sabijun.github.io/MT_RewardTreePage.
@@
Zhaopeng Feng1♠  Jiahan Ren1♠   Jiayuan Su1♠   Jiamei Zheng1 Hongwei Wang1†   Zuozhu Liu1† 1Zhejiang University {zhaopeng.23,jiahan.24,jiayuan.23,jiamei.24}@intl.zju.edu.cn {hongweiwang,zuozhuliu}@intl.zju.edu.cn
@@ 1 Introduction
Machine translation (MT) naturally aligns with token-level MDP frameworks, as each translation decision corresponds directly to token generation. However, there is still a lack of systematic methodologies for constructing and evaluating PRMs in MT, which has hindered progress relative to advancements in general-domain LLMs.
@@
Developing effective PRMs is challenging. Although Lightman et al. (2024) demonstrate that process supervision with human annotators improves PRM performance in mathematical tasks, this methodology requires domain-expert annotators, resulting in prohibitive costs and practical limitations for translation tasks. Recently, some studies suggest that a PRM can be automatically learned during Direct Preference Optimization (DPO) training (Rafailov et al., 2024b, a; Yuan et al., 2024). However, existing vanilla preference pair datasets provide only sequence-to-sequence preference data, rather than token-level preferences, which raises concerns about their applicability for token-level alignment. Additionally, evaluating PRMs remains a significant challenge. In mathematical tasks, evaluation is often done using a Best-of-N (BoN) sampling strategy—selecting the highest-scored response from N candidates based on a PRM (Lightman et al., 2024; Wang et al., 2024b; Luo et al., 2024)—or by having the PRM identify errors or verify correctness in the steps (Zheng et al., 2024; Zhang et al., 2025). Since each step in mathematics has a deterministic answer, these methods do not directly translate to PRM evaluation in MT.
@@
In this paper, we introduce MT-RewardTree, a comprehensive framework for constructing, evaluating, and deploying PRMs in machine translation. We propose an approximate Monte Carlo Tree Search (MCTS) method (Kocsis and Szepesvári, 2006; Silver et al., 2016) to generate the token-level preference pair dataset. This dataset is then split into a training set for reward model development and a benchmark for reward evaluation. We provide a systematic comparison of different reward modeling methods and test on both token-level and sequence-level performance. Furthermore, we demonstrate two practical applications of PRMs, offering valuable insights for future MT research. Our main contributions are as follows:
  • We introduce MT-RewardTree, a comprehensive framework for the construction, evaluation, and deployment of PRMs in MT. We establish the first dedicated reward benchmark - MT-PRMBench. Our experiments demonstrate that MT-PRMs achieve competitive performance on both token-level and sequence-level evaluations.
@@
  • Comprehensive experiments indicate that our token-level preference pairs, generated through an approximate MCTS method, significantly outperform vanilla preference pairs in process reward model training. Furthermore, our analysis validates that supervising PRMs using preference-based signals is more effective than direct supervision with absolute value estimates.
  • We demonstrate that our MT-PRMs can directly identify token-level translation differences and facilitate test-time alignment for LLM-based MT without the need for additional alignment training, offering valuable practical insights for the application of reward models in MT.
@@ 2.1 Token-level Markov Decision Process
where $(x_{0},\dots,x_{L})$ represents the input prompt and $(y_{0},\dots,y_{t-1})$ is the sequence of generated tokens until time step t1t-1. The state transition function $f$ is deterministic and updates the state by concatenating the newly generated token:
@@ 2.2 Reward Modeling in RLHF
Although token-level reward modeling offers finer-grained feedback, obtaining effective PRMs is more challenging to obtain and deploy (Lightman et al., 2024; Cao et al., 2024).
@@ 3.1 Constructing Token-level Preference Pairs
Drawing inspiration from MCTS, we propose a token-centric approach that quantifies token quality based on its potential to contribute to higher-quality translations. This method aligns with Monte Carlo-based PRMs construction techniques in mathematics, where step-wise quality is determined by its incremental contribution to deriving correct answers (Wang et al., 2024b; Guan et al., 2025).
@@
Model MT-PRMBench
Sequence-level Token-level
EN→$X$ XX→EN Avg. EN→$X$ XX→EN Avg.
@@
Skywork-Reward-LLaMA-3.1-8B 0.857 0.773 0.815 - - -
MT-Ranker-base 0.785 0.787 0.786 - - -
MT-Ranker-large 0.847 0.873 0.860 - - -
[.4pt/2pt]      PRMs
MT-PRM-LLaMA-3.2-3B 0.777 0.775 0.776 0.542 0.615 0.578
MT-PRM-Qwen-2.5-3B 0.867 0.858 0.863 0.637 0.685 0.660
Table 1: Accuracy results on MT-PRMBench. Skywork-Reward-LLaMA-3.1-8B is an advanced ORM for general domains, while MT-Ranker represents the SoTA non-metric reference-free translation quality estimation model.
Translation Direction Token-level Preference Pairs
Train MT-PRMBench
DE-EN 1,255 200
EN-DE 2,059 200
RU-EN 1,219 200
EN-RU 1,711 200
ZH-EN 1,232 200
EN-ZH 1,176 200
Table 2: Data Statistics.
@@ 3.2 Implicit Process Reward Modeling
Preference Pair Type Training Strategy Avg.
Token-level DPO 0.660
Vanilla DPO 0.574
Token-level KTO 0.644
Vanilla KTO 0.562
Table 3: Ablation study on the effect of training preference data and implicit reward training objectives. The backbone model is Qwen-2.5-3B-Instruct and we test these variants on MT-PRMBench (Token-level).
4 Experiments
4.1 Experimental Setup
Datasets. We explore four languages—English (EN), German (DE), Chinese (ZH), and Russian (RU)—and six translation directions: EN→XX and XX→EN. Our raw corpus consists of test sets from WMT17 to WMT20, supplemented with development and test sets from the Flores (Costa-jussà et al., 2022). We use the TowerInstruct-7B-v0.2111https://huggingface.co/Unbabel/TowerInstruct-7B-v0.2 model with a temperature of 0.95 and apply the MCTS-based approach described earlier. During the Simulation step, we sample three candidate hypotheses for each node. Our token-level preference pairs is divided into train and test set (MT-PRMBench). Detailed statistics are in Table 2.
Training Details. We take LLaMA-3.2-3B-Instruct and Qwen-2.5-3B-Instruct as the backbone models for training. For the DPO training, the higher-scored sentence is designated as the chosen response, while the lower-scored sentence is labeled as the rejected response. For the KTO training, the higher-scored sentence is treated as the positive sample, and the lower-scored sentence as the negative sample. We set $\beta$ as 0.1.
@@ 4.1 Experimental Setup
MT-PRMBench. MT-PRMBench comprises two distinct subsets for evaluation: Token-level and Sequence-level. The Token-level subset is designed for assessing preferences between immediate next-token candidates that follow an identical input prefix. In contrast, the Sequence-level facilitates the comparison of entire generated sequence completions that also originate from a shared input prefix.
4.2 Evaluation Results
@@
Token-level Performance. From Table 3.1, we can observe that our MT-PRM-LLaMA-3.2-3B and MT-PRM-Qwen-2.5-3B models achieved accuracies of 0.578 and 0.66 respectively on the token-level MT-PRMBench. As shown in Table 3, we systematically compare models trained with vanilla sequence-level preference pairs versus our token-level preference pairs, while evaluating both DPO and KTO training objectives. The results demonstrate that token-level preference pairs significantly improve discrimination accuracy: implicit PRMs trained with token-level preference pairs outperform vanilla sequence-level baselines by +8.6% (DPO) and +11.5% (KTO). This performance gap highlights the critical advantage of token-level preference pairs in helping capture fine-grained translation quality distinctions.
@@
Sequence-level Performance. We also convert our PRMs to sequence-level scoring through weighted DPO rewards (as shown in Eq. 11). We can observe that our MT-PRM-Qwen-2.5-3B achieves the highest performance among all models in the Prefixed set, with an average score of 0.863, outperforming both Skywork-Reward-LLaMA-3.1-8B  [^2] https://huggingface.co/Skywork/Skywork-Reward-Llama-3.1-8B and the MT-Ranker (Moosa et al., 2024) variants. This demonstrates the effectiveness of our token-level supervision framework even when adapted to sequence-level scoring.
MT-PRMBench (Token-level) DE-EN RU-EN ZH-EN EN-DE EN-RU EN-ZH Avg.
Supervised by Value (SV) 0.52 0.56 0.48 0.50 0.51 0.55 0.52
Supervised by Preference (SP) 0.64 0.74 0.68 0.58 0.68 0.66 0.66
Table 4: Experimental results of two training methods on the Token-level Benchmark
Source 在橘红色背景墙映衬下格外鲜亮的广告牌上,它们各自标出了不菲的价码。
[.4pt/2pt]      Reference On the bright advertising board against a tangerine colored wall, they listed their respective exorbitant price tags.
[.4pt/2pt]      Translation A Against an orange background, the bright billboards listed their respective, hefty prices.
Translation B Against an orange background, the dark billboards listed their respective, hefty prices.
Translation C Against an orange background, the bright billboards listed their respective, bargain prices.
Translation A ’Against’ ’ an’ ’ the’ ’ bright’ ’ bill’ ’ hefty’ ’ prices’ Weighted Implicit Rewards(↑) COMETKiwi(↑)
Reward -2.3 -2.3 0.07 1.28 -2.3 0.03 0.09 -3.43 0.80
[.4pt/2pt]      Translation B ’Against’ ’ an’ ’ the’ ’ dark’ ’ bill’ ’ hefty’ ’ prices’ Weighted Implicit Rewards(↑) COMETKiwi(↑)
Reward -2.3 -2.3 0.07 -2.3 -2.3 0.09 0.09 -4.13 0.73
[.4pt/2pt]      Translation C ’Against’ ’ an’ ’ the’ ’ bright’ ’ bill’ ’ bargain’ ’ prices’ Weighted Implicit Rewards(↑) COMETKiwi(↑)
Reward -2.3 -2.3 0.07 1.28 -2.3 -0.95 -2.3 -3.99 0.78
Table 5: Case study illustrating token-level credit assignment by our Qwen-PRM.
5 Analysis and Practical Insight
5.1 Modeling Advantage versus Value as the PRM Training Signal
We have explored the impact of vanilla preference pairs versus MCTS-generated token-level preference pairs on the performance of implicit PRMs in the previous experiments (Table 3). This section shifts focus to the nature of the supervisory signal used for training PRMs. Specifically, we analyze our choice of implicit reward modeling—derived from token-level preference pairs (termed "Supervised by Preference (SP)")—against the alternative of directly using MCTS-backpropagated values $V(a)$ as a dense supervisory signal (termed "Supervised by Value (SV)").
Table 4 demonstrates that SP yields markedly superior performance in token-level discrimination, where SP achieved an average accuracy of 0.66, while SV achieved 0.52. This significantly higher accuracy underscores the effectiveness of our implicit PRM when trained with preference-based supervision. The SV approach, directly regressing on MCTS-backpropagated $V(a)$ values, tasks the PRM with learning a value function. However, $V(a)$ as a dense, token-level reward faces limitations: Monte Carlo estimates can be noisy, and the absolute value of a partial translation may offer an indirect and insufficiently discriminative signal for the most recent token’s quality, especially in complex MT scenarios or with imperfect rollouts—a recognized challenge in process supervision literature (Lightman et al., 2024). We also find that these $V(a)$ values often clustered within a narrow numerical range (e.g., 0.7-0.8) further illustrates why this SV approach struggles. Such clustering severely hampers a regression model’s ability to discern fine-grained distinctions, as the supervisory signal becomes subtle.
In contrast, the SP approach, using DPO/KTO objectives, excels by directly learning a representation reflecting the advantage of one token choice over another. This implicitly shapes a reward function where the per-step process reward, $r_{\theta}^{t}=\beta\log\frac{\pi_{\theta}(y_{t}|y_{<t})}{\pi_{ref}(y_{t}|y_{<t})}$ (Eq. 9), quantifies the localized, step-wise advantage of selecting token $y_{t}$. This preference-based optimization strategy is well-supported by prior work (Yuan et al., 2024; Rafailov et al., 2024a; Wang et al., 2024a). Crucially, DPO’s mechanism, focusing on the log-probability ratio between sequences (Eq. 4), is inherently more sensitive to relative differences than absolute value regression. This makes it adept at capturing preferences even when underlying $V(a)$ scores from the SV context are close, explaining its superior performance in training PRMs for nuanced token-level discrimination in our MT setting.
5.2 Per-token Credit Assignment
Our PRM can identify token-level translation differences, with the per-step process reward defined as Eq. 9 quantifying the reward for generating token $y_{t}$ at step $t$. These individual token rewards are then aggregated into a final weighted sequence score (Eq. 11), allowing for a comprehensive evaluation that originates from fine-grained assessments.
Table 4.2 provides a case study illustrating these capabilities. For instance, in Translation B, the token "dark", which semantically contradicts the source " 鲜亮 " (bright), receives a significantly negative reward (-2.3) from our PRM. In Translation C, the incorrect token "bargain", used where " 不菲的 " (exorbitant/costly) is implied, is substantially penalized (-0.95). In contrast, contextually appropriate tokens in Translation A, such as "bright" (1.28) and "hefty" (0.03), secure relatively higher rewards. Furthermore, the final weighted sequence scores computed by our PRM demonstrate strong alignment with automatic metrics like COMETKiwi. Translation A, the highest quality hypothesis (COMETKiwi 0.80), also achieves the most favorable weighted PRM score (-3.43). This case study thus substantiates our PRM’s effective token-level credit assignment and the consistency of its fine-grained assessments with established sequence-level quality metrics.